2. A patient receives 3.5 L of glucose solution intravenously (IV). If 100. mL of the solution contains 5.7 g of glucose (carbohydrate), how many kilocalories did the patient obtain from the glucose solution? (Express the energy to two significant figures and include the appropriate units.)
Answer #1

total amount of glucose obtained by patient = 5.7*3.5*10^3/100

                          = 199.5 g

no of mol of glucose obtained = 199.5/180 = 1.11 mol

enthalpy of combustion of glucose(DHcomb) = -2800 kj/mol

amount of energy obtained by patient = 2800*1.11 = 3108 kj

                                       = 742.83 kcal

                                      = 7.4*10^2 kcal